Two bodies of masses m and 4 m are placed at a distance r. The gravitational potential at a point on the line joining them where the gravitational field is zero is:

Grade:12

2 Answers

Manas Satish Bedmutha
9 years ago
Field is GM / R2, in general.
So, at  x distance from m, field = Gm/x2,
And for 4m, distance r-x, field is G (4m) / (r-x)2
Equating both, r-x / x = 2. that is at r/3 from m
 
aatif ahmad
5 years ago
For M,gfi=GM/x^2 similarly for 4M gfi=4GM/(r-x)^2..dividing both eq.ns we get r-x/x=2 therefore,x=r/3 now g.p= -G(4M-3M)/x ,we get g.p= -9GM/r